Solve the following initial value problem d^2 y/dx^2 +dy/dx -2y =-6 sin 2x -18 cos 2x , y(0) = 2, y'(0) = 2
1
Expert's answer
2020-05-22T18:19:05-0400

The given equation can be written as

(D2+D-2)y = -6 sin2x -18cos2x


The auxiliary equation of the above is

D2 +D-2=0

D = 1,-2

The complimentary solution is given by:

y= c1 ex + c2 e-2x


Particular solution is given by:

"\frac{1}{D^2+D-2}( -6sin2x -18cos2x )="

"=\frac{1}{D^2+D-2} (-6sin2x)-\frac{1}{D^2+D-2}(-18cos2x)"


"=\frac{-6sin2x}{D-6} - \frac{18cos2x}{D-6}"


"= (D+6) \frac{-6sin2x}{D^2-36}-(D+6) \frac{18cos2x}{D^2-36}"


"=(D+6) \frac{-6sin2x}{-40} -(D+6) \frac{18cos2x}{-40}"


"=\frac{3}{10}(cos2x+3sin2x) + \frac{9}{10} (-sin2x+3cos2x)"

=3 cos2x


So, the solution is y= c1 ex+ c2e-2x + 3 cos2x


Now, initial conditions given are:

y(0)=2

-1=c1+c2

y' = c1ex -2c2e-2x -6 sin2x

y'(0)=2

2= c1-2c2

c1=0 , c2=-1


y=-e-2x +3cos2x